OpenStudy (gabylovesyou):

Which equation below is not an equation that represents the graph shown below?

OpenStudy (gabylovesyou):

y + 2 = -3(x - 1) 3x + y =1 y - 1 = -3(x + 2) y = -3x + 1

OpenStudy (anonymous):

what is the y intersect of this line ? what is the slope of this line ?

OpenStudy (gabylovesyou):

the slope is 1/3

OpenStudy (gabylovesyou):

idk bout the y intercept

OpenStudy (anonymous):

why do you say that the slope is 1/3 ? first the slope should be negative since the graph is decreasing .. next y intercept is the point where the graph touches the y axis

OpenStudy (gabylovesyou):

sory 1/-3

OpenStudy (anonymous):

it's -3 the slope is delta y / delta x or you can think of it : for every unit that you go towards the positive x direction how much does the y changes ? as you can see here for every unit in the x direction we get down by 3 in the y direction

OpenStudy (gabylovesyou):

ohh ok... what is the y intercept

OpenStudy (anonymous):

where does the graph touches the y axis ? at y = 1 so in the slope intercept form : y = ax + b a is the slope and b is the y intercept so our graph equation should be y = -3x + 1

OpenStudy (anonymous):

no you have to check which one of the equations isnt like this

OpenStudy (gabylovesyou):

so the answer would be D?

OpenStudy (anonymous):

we just said that y = -3x + 1 describes the graph

OpenStudy (anonymous):

your question is which of the equation is NOT representing the graph

OpenStudy (gabylovesyou):

ohhh ok.....

OpenStudy (anonymous):

so what is the answer?

OpenStudy (gabylovesyou):

it has to be between a and c right?

OpenStudy (anonymous):

yes since we said that it isnt D

OpenStudy (anonymous):

you just have to simplify the equations and see which one of them doesnt look like y=-3x+1

OpenStudy (anonymous):

i have to go now so ill give the answer with explanation : A.y + 2 = -3(x - 1) y +2 = -3x + 3 y = -3x + 1 not the answer B. 3x + y =1 y = -3x +1 not the answer C. y - 1 = -3(x + 2) y -1 = -3x - 6 y = -3x - 5 this is the answer

OpenStudy (gabylovesyou):

thank u!